9lbs of mixed nuts containing 55% peanuts were mixed with 6lbs of another kind of mixed nuts that contain 40% peanuts. What percent of the new mixture is peanuts?
Answer by checkley75(3666)

.55*9+.40*6=15x
4.95+2.4=15x
7.35=15x
x=7.35/15
x=.49 or 49% of the new mix is peanuts.